OpenStudy (anonymous):

What is this complex fraction simplified? 3/(2/x + y) Please show the steps as well. Thanks! :)

OpenStudy (amonoconnor):

You bet! Hold on second...

OpenStudy (amonoconnor):

Is the 2 attached to the x? Is it 3/((2/x) + y) Or 3/(2/(x + y))

OpenStudy (amonoconnor):

I see where I went wrong... I'm working on it :)

OpenStudy (anonymous):

Thanks a bunch! :)

OpenStudy (amonoconnor):

Hey! Sorry if I took too long- I got sidetracked with something else, but. I totally have what you're asking for! Here you go: 3/((2/x) + y) = 3/((2/x) + (xy/x)) = 3/((2 + xy)/x) = 3(x/(2 + xy)) = 3x/(2 + xy) There you go! :)

OpenStudy (amonoconnor):

@Unknown11

OpenStudy (amonoconnor):

Does that help? Does that make sense? :)

OpenStudy (anonymous):

Thanks a lot for your help! Makes a lot more sense now @amonoconnor. :)